# Wavy Transformer

### v1

@@ 5.1 Experiments: NLP Tasks
5.1.1 Experimental Settings
Here, we adopt a smaller-scale pretraining setup than standard BERT configurations to facilitate faster experimentation. We implement three types of models with different residual connections such as diffuse, wave, and mix in a 24-layer BERT with hidden size 256 and 4 attention heads based on Post-LN Wavy Transformer blocks. As for feed-forward layer, we set the intermediate size to 1024. For both wave and mixed residual connections, we set the time interval to $\tau=0.5$ unless stated otherwise. In this experiment, the mixing coefficient was defined as $\lambda=\textrm{sigmoid}(\theta)$, with $\theta$ being a single scalar parameter initialized to 0; consequently, the same $\lambda$ is shared across all feature dimensions. For the mixing strategy, we adopt the velocity-based variant: $\mathbf{Y}^{l+1}=\lambda\bigl{[}\tau(\mathbf{A}-\mathbf{I})\mathbf{X}^{l}+\mathbf{Y}^{l}\bigr{]}+(1-\lambda)(\mathbf{A}-\mathbf{I})\mathbf{X}^{l}.$ As a optimizer, we use AdamW loshchilov2018decoupled with learning rate $5.0\times 10^{-5}$, warm-up for first 10 % of steps, then linear decay to step 10k. Also, we evaluate every 5k steps on the validation set, reporting perplexity (PLP) and MLM accuracy. Also, the other hyper-parameters for different downstream tasks are given in Appendix G.
5.1.2 Perplexity and MLM Accuracy
First, we present the pretraining results to illustrate the fundamental capacity of each model with its respective residual connection as a language model. Table 1 summarizes the validation perplexity and MLM accuracy after 10k steps of batch size 64 for each model.
Interestingly, while the wavy residual connection alone just achieves nearly the same-level performance as the diffusive residual connection, the mixed residual connection outperforms both. This suggests that the wavy residual connection capture fundamentally different or complementary dynamics compared to the usual diffusive residual connection, and that gating both types of updates together with a trainable $\lambda$ enables the model to leverage a broader range of interactions than a purely diffusive attention update.
Table 1: Validation perplexity (PPL) and MLM accuracy (MLM Acc) after 10k steps. Arrow symbols denote that lower PPL and higher MLM Acc are better.
@@ 5.1.2 Perplexity and MLM Accuracy
Diffuse 31.76 44.39 %
Wave 31.99 44.52 %
Mix 29.00 45.56 %
@@ 5.1.3 GLUE Fine-tuning Performance
The GLUE wang2018glue benchmark provides a comprehensive suite of natural language understanding tasks such as sentence similarity (STS-B), paraphrase detection (MRPC, QQP), and textual entailment (RTE) that serve as critical tests for evaluating the performance of pretrained language models. In our experiment, we fine-tuned each pretrained BERT-like model on each GLUE task dataset for demonstrating advantages of Wavy Transformer over baseline methods.
Table 2 reports the mean GLUE scores (across three random seeds) for each task, as achieved by BERT models employing various residual connections. The mixed (Diffusion + Wave) residual model achieves the highest overall GLUE score, driving most of its gains through stronger performance on CoLA, SST-2, RTE and WNLI compared to the pure Diffuse or Wave variants. Although it dips slightly on STS-B, the mixed approach clearly balances improvements across both classification and inference tasks. Together with Table 1, these results strongly suggest (i) that Wavy Transformer can capture different features unattainable by conventional diffusive connections, and (ii) that combining these different types of features can enhance BERT’s performance on NLP tasks.
@@
Diffuse 10.67 83.64 76.36 83.96 73.51/73.03 81.63 52.95 51.11 14.10 67.00
Wave 10.64 83.10 76.11 84.17 72.10/72.94 80.74 53.67 56.34 11.65 66.83
Mix 12.07 84.25 76.32 83.81 73.04/73.53 81.19 55.35 55.40 12.46 67.49
@@ 5.1.4 Analysis of Over-smoothing
Figure 3: Cosine similarity across layers for a mix connection with 1-sigma interval shading.
We further conduct the analysis of the over-smoothing behavior for each residual connections. Especially, we computed the cosine similarity as a measure to investigate the over-smoothing behavior following shi2022revisiting : $\mathrm{CosSim}=\frac{1}{n(n-1)}\sum_{i\neq j}\frac{\mathbf{x}_{i}^{\top}\mathbf{x}_{j}}{\|\mathbf{x}_{i}\|\,\|\mathbf{x}_{j}\|}.$
For computation of the cosine similarity, we use data randomly sampled from WIkipedia and BookCorpus as input to the BERT-based pretraining models fine-tuned on the SQuAD dataset rajpurkar-etal-2016-squad .
Figures 3 and 3 present a comparison of cosine similarity across layers. The Figure 3 illustrates the results for diffusive and wavy residual connections, while the Figure 3 shows those for mixed residual connections with 1-sigma interval shading. As shown in Figure 3, under the current conditions, the wave and diffuse residual connections did not exhibit any significant differences in terms of the degree of over-smoothing. In other words, the values of cosine similarity including their variances are almost same at the 24-th layers for diffuse and wave residual connections. However, the behaviors were markedly different. In the case of the diffusive residual connections, a monotonous increase in cosine similarity is observed as the layer depth increases. In contrast, the wavy residual connections exhibit oscillatory behavior in cosine similarity. Also, it decreases near the final layer, which is markedly different from the diffusive case where similarity monotonically approaches a uniform state. This result appropriately reflects the update rule based on the wave equation. Consequently, the results suggest that diffusive and wavy residual connections extract fundamentally different types of features from the data, which may be related to the frequency of hidden states across layers.
In Figure 3, the behavior of the hidden states for the mixed residual connections is shown. Taken together with the results in Tables 1 and 2, this indicates that the features extracted by the wavy residual connection complement those derived from the standard diffusive residual connection, thereby enhancing the performance of the baselines by broadening the range of extracted features.
@@ 5.2 Experiments: CV Tasks
In this section, we demonstrate the capacity of the proposed approach for CV tasks. All of our experiments are conducted on the ImageNet dataset imagenet_cvpr09 including around 1k classes 1.3M images in the training dataset and 50k images in the validation set.
5.2.1 Experiment Settings
To demonstrate that our proposed approach are beneficial to CV tasks, we choose 12-layer DeiT DeiT as our backbone. Also, for evaluating its consistency with the other remedy to avoid over-smoothing, we combine our approach with existing technique such as FeatScale wang2022antioversmooth . When we use 12-layer DeiT, we follow the same training recipe, hyper-parameters, and data augmentation with DeiT ; wang2022antioversmooth . We implemented DeiT based on Pre-LN Wavy Transformer blocks considering that original DeiT is based on Pre-LN Transformer block. We set $\tau=0.5$, and we select $\mathbf{X}^{l+1}=\boldsymbol{\lambda}\,\mathbf{X}_{\text{wave}}^{l+1}+(1-\boldsymbol{\lambda})\,\mathbf{X}_{\text{diffuse}}^{l+1},$ as the mixing strategy and initialize the gating parameter $\boldsymbol{\lambda}$} as an all-zero vector whose length equals the dimension of the hidden states in the mix variant.
@@ 5.2.2 Results of ImageNet Classification
Table 3 summarizes our experimental results. The top two rows are baseline scores taken from wang2022antioversmooth . Rows labeled “Diffuse+Wave” in the residual column are our proposed extensions. These results demonstrate the effectiveness and generality of Wavy Transformer: with almost ignorable minimal additional parameters, it also consistently improves baseline models in CV tasks. Also, an analysis of over-smoothing for the ImageNet classification task is provided in Appendix D.
Table 3: Experimental evaluation of Wavy Transformer block plugged into DeiT DeiT . The number inside the ($\uparrow\cdot$) represents the performance gain compared with the model without Wavy Transformer block. The top two rows are baseline results taken from nguyen2023mitigating .
@@
DeiT-Ti Diffuse 224 12 5.7M 72.17
DeiT-Ti + FeatScale Diffuse 224 12 5.7M 72.35
DeiT-Ti Diffuse + Wave 224 12 5.7M 72.33  ($\uparrow\,$0.16)
DeiT-Ti + FeatScale Diffuse + Wave 224 12 5.7M 72.62  ($\uparrow\,$0.26)
@@ 6 Conclusions
In this paper, we investigated the dynamics embedded in attention layers from the viewpoint of graph neural diffusion and introduced Wavy Transformer. We first established an equivalence between the intrinsic dynamics driven by attention layers and a graph neural diffusion process on a complete graph. From this perspective, the over-smoothing behavior can be understood as the dissipative nature of diffusion. Based on this insight, we proposed Wavy Transformer, which can be seamlessly integrated into standard diffusive transformer architecture. Our extensive experiments on NLP and CV tasks showed that it can capture feature dynamics that is qualitatively different from those driven by conventional transformers, and that integrating it into existing models enhances performance without increasing the number of training parameters or requiring additional hyperparameter tuning. We believe this work represents a first step toward a physics-inspired design of transformer architecture.
Appendix A Energy Evolution in Diffusion and Wave Systems

### v2

@@ 5.1 Experiments: NLP Tasks
Here, we adopt a smaller-scale pretraining setup than standard BERT configurations to facilitate faster experimentation. The implementation details are given in Appendix A.1.1
5.1.1 PPL and MLM Accuracy
First, we present the pretraining results to illustrate the fundamental capacity of each model with its respective residual connection as a language model. Table 1 summarizes the validation PPL and MLM accuracy after 10k steps of batch size 64 for each model. Interestingly, while the wavy residual connection alone just achieves nearly the same-level performance as the diffusive residual connection, the mixed residual connections outperform both. This suggests that the wavy residual connection capture fundamentally different or complementary dynamics compared to the usual diffusive residual connection, and that gating both types of updates together with a trainable $\boldsymbol{\lambda}$ enables the model to leverage a broader range of interactions than a purely diffusive attention update.
Table 1: Validation PPL and MLM accuracy (MLM Acc) after 10k steps. Arrow symbols denote that lower PPL and higher MLM Acc are better.
@@ 5.1.1 PPL and MLM Accuracy
Diffusion 31.76 44.39 %
Full Wave 31.99 44.52 %
Mix (+Full) 29.00 45.56 %
Mix (+Light) 32.29 44.53 %
@@ 5.1.2 GLUE Fine-tuning Performance
The GLUE wang2018glue benchmark provides a comprehensive suite of natural language understanding tasks that serve as critical tests for evaluating the performance of pretrained language models. In our experiment, we fine-tuned each pretrained BERT-like model on each GLUE task dataset for demonstrating advantages of Wavy Transformer over baseline methods. As summarized in Table 2, the mixed residual with the momentum-only variant (Mix with Light Wave) attains the best macro average over the 10 metrics (+1.99+1.99 over Diffusion). Gains are concentrated on CoLA, QNLI, RTE, WNLI, and especially STS-B (+12.65+12.65 vs. Diffusion), while remaining competitive on MNLI-m; minor regressions appear on SST-2, MRPC, QQP, and MNLI-mm. The Mix with Full Wave variant reaches the top SST-2 but underperforms overall due to a large STS-B drop with the corrected evaluator. Together with Table 1, these results strongly suggest (i) that Wavy Transformer can capture different features unattainable by conventional diffusive connections, and (ii) that combining these different types of features can enhance BERT’s performance on NLP tasks.
@@
Diffusion 10.67 83.64 76.36 83.96 73.51/74.03 81.63 52.95 51.11 52.11 64.13
Full Wave 10.64 83.10 76.11 84.17 72.10/72.94 80.74 53.67 56.34 32.91 62.27
Mix (+Full) 12.07 84.25 76.32 83.81 73.04/73.53 81.19 55.35 55.40 29.40 62.44
Mix (+Light) 12.77 82.76 76.18 83.89 73.71/73.77 81.98 55.96 55.40 64.76 66.12
@@ 5.1.3 Analysis of Over-smoothing
Figure 3: Cosine similarity across layers for a mix (+ Full Wave) with 1-sigma interval shading.
We further conduct the analysis of the over-smoothing behavior for each residual connections. Especially, we computed the cosine similarity as a measure to investigate the over-smoothing behavior following shi2022revisiting : $\mathrm{CosSim}=\frac{1}{n(n-1)}\sum_{i\neq j}\frac{\mathbf{x}_{i}^{\top}\mathbf{x}_{j}}{\|\mathbf{x}_{i}\|\,\|\mathbf{x}_{j}\|}.$ For computation of the cosine similarity, we use data randomly sampled from WIkipedia and BookCorpus as input to the BERT-based pretraining models fine-tuned on the SQuAD dataset rajpurkar-etal-2016-squad . Figures 33 compare layerwise cosine similarity: Figure 3 contrasts the difference between diffusive and wavy, and Figure 3 shows the mixed residual with Full Wave (with $1\sigma$ shading). As seen in Figure 3, diffusive and wavy residuals exhibit similar over-smoothing at the 24th layer (means and variances are comparable), but their trajectories markedly differ: diffusion increases monotonically with depth, whereas wave oscillates and drops near the final layer, consistent with wave-equation updates. This suggests that diffusive and wavy residual connections extract fundamentally different types of features from the data, which may be related to the frequency of hidden states across layers. Figure 3 shows the mixed residual; together with Tables 1 and 2, this supports that wavy features complement diffusive ones, and that gating them broadens the feature range and improves performance.
@@ 5.2 Experiments: CV Tasks
In this section, we demonstrate the capacity of the proposed approach for CV tasks. All of our experiments are conducted on the ImageNet dataset imagenet_cvpr09 including around 1k classes 1.3M images in the training dataset and 50k images in the validation set. Also, we choose 12-layer DeiT DeiT as our backbone. The implementation details is given in Appendix A.2.1.
@@ 5.2.1 Results of ImageNet Classification
Table 3 summarizes our experimental results. The two rows labeled "Diffusion" in the residual column are baseline scores taken from nguyen2023mitigating . The other rows with "+ Wave" are our proposed extensions. These results demonstrate the effectiveness and generality of Wavy Transformer: with almost ignorable minimal additional parameters, it also consistently improves baseline models in CV tasks.
Table 3: Experimental evaluation of Wavy Transformer block plugged into DeiT DeiT . The number inside the ($\uparrow\cdot$) represents the performance gain compared with the model without Wavy Transformer block. The two rows labeled "Diffusion" are baseline results taken from nguyen2023mitigating .
@@
DeiT-Ti Diffusion 224 12 5.7M 72.17
DeiT-Ti Diffusion + Full Wave 224 12 5.7M 72.33  ($\uparrow\,$0.16)
DeiT-Ti Diffusion + Light Wave 224 12 5.7M 73.09  ($\uparrow\,$0.92)
DeiT-Ti + FeatScale Diffusion 224 12 5.7M 72.35
DeiT-Ti + FeatScale Diffusion + Full Wave 224 12 5.7M 72.6272.62\, ($\uparrow\,$0.26)
Quantifying Over-Smoothing. Beyond cosine similarity, we quantify the spectral gap of the attention matrix and node-feature variance. Wave dynamics show a smaller gap and larger variance on CV tasks (Table 4); definitions and additional results on trends of cosine similarity are in Appendix A.2.2.
Table 4: Over-smoothing diagnostics on CV tasks. Values are calculated over validation images. Spectral gap is $1-|\lambda_{2}(\textbf{A})|$ for the attention matrix. Lower gap ($\downarrow$) and higher node ($\uparrow$) indicate mitigated over-smoothing and Inter-class variance is reported only as an auxiliary separability measure.
Dynamics Spectral gap $(1-|\lambda_{2}|)$ Node-feature var. Inter-class var.
Diffusion 0.836 $\pm$ 0.00345 2.480 $\pm$ 0.0778 0.195
+ Full Wave 0.629 $\pm$ 0.00887 2.609 $\pm$ 0.0902 0.211
+ Light Wave 0.730 $\pm$ 0.00842 2.109 $\pm$ 0.0696 0.308
5.2.2 Scaling to Deeper Models
We additionally evaluate CaiT-XXS-24 on ImageNet-1K under the same training recipe as the baseline Touvron_2021_ICCV . The wave residual improves Top-1 accuracy by $1.0$ point. The 0.90.9M parameter reduction arises solely from omitting extra class-attention block; all other settings remain unchanged (Table 5).
Table 5: ImageNet-1K (CaiT-XXS-24). Numbers in $(\uparrow\cdot)$ indicate Top-1 gain over the diffusion-only baseline. The training recipe and input resolution are identical to the baseline Touvron_2021_ICCV .
Method Residual Input Size #Layer #Param Top-1 Acc (%)
CaiT-XXS-24 Diffusion 224 24 12.0M 77.6
Wavy Transformer Diffuse + Full Wave 224 24 11.1M 78.6(1.0)\,($\uparrow$ 1.0)
5.3 Experiments: Sparse-Graph Tasks
We further evaluate on OGBN-Arxiv and OGBN-Proteins hu2020ogb using DIFFormer wu2023difformer under the authors’ training protocol and hyperparameters. Diffusion denotes vanilla DIFFormer; + Light Wave denotes addition of our momentum-only wave residual. Table 6 highlights representative depths. Complete results are provided in Appendix A.3.2. + Light Wave substantially mitigates depth collapse on OGBN-Arxiv and delivers a consistent ROC-AUC gain on OGBN-Proteins.
Table 6: OGB highlights (mean $\pm$ standard deviation over 3 seeds). Arxiv reports accuracy and Proteins reports ROC-AUC; all values are percentages. $\Delta$ denotes absolute improvement.
Dataset Metric $\uparrow$ #Layer Diffusion + Light Wave $\Delta$
OGBN-Arxiv Acc 7 24.44 $\pm$ 4.51 66.73 $\pm$ 0.33 +42.29
OGBN-Proteins ROC-AUC 5 69.42 $\pm$ 2.31 80.14 $\pm$ 0.67 +10.72
5.4 Computational Cost
For cost-sensitive settings, we adopt the momentum-only variant (Light Wave), which preserves the second-order residual with negligible overhead. On both BERT and DeiT-Tiny, throughput and memory stay within a few percent of the diffusion baseline (Table 7); complete cost tables, including the Full Wave variant, are provided in Appendix A.4.
Table 7: Computational efficiency of each residual connections on 4 Nvidia V100.
Model Variant Inference Training Peak GPU Mem
BERT Diffusion 101.6 415.6 18.31
BERT Light Wave 101.3 436.2 18.69
DeiT-Tiny Diffusion 2631.1 618.6 8.25
DeiT-Tiny Light Wave 2644.2 617.6 9.14
@@ 6 Conclusions
In this paper, we investigated the dynamics embedded in attention layers from the viewpoint of graph neural diffusion and introduced Wavy Transformer. We first established an equivalence between the intrinsic dynamics driven by attention layers and a graph neural diffusion on a complete graph. From this perspective, the over‑smoothing behavior can be understood as the dissipative nature of diffusion. Based on this insight, we proposed Wavy Transformer, which can be seamlessly integrated into standard diffusive transformer architecture. Our extensive experiments on NLP, CV, and sparse-graph tasks showed that it can capture feature dynamics that is qualitatively different from those driven by conventional transformers, and that integrating it into existing models enhances performance without increasing the number of training parameters or requiring additional hyperparameter tuning. We believe this work represents a first step toward a physics‑inspired design of transformer architecture.
Acknowledgement
A part of numerical simulations were conducted using Earth Simulator at JAMSTEC.
Appendix A Experimental Details and Additional Results
A.1 NLP Tasks
A.1.1 Implementation Details
We implement three types of models with different residual connections such as diffuse, wave, and mix in a 24-layer BERT with hidden size 256 and 4 attention heads based on Post-LN Wavy Transformer blocks. As for feed-forward layer, we set the intermediate size to 1024. For both wave and mixed residual connections, we set the time interval to $\tau=0.5$ unless stated otherwise. In this experiment, the mixing coefficient was defined as $\lambda=\textrm{sigmoid}(\theta)$, with $\theta$ being a single scalar parameter initialized to 0; consequently, the same $\lambda$ is shared across all feature dimensions. For the mixing strategy with Full Wave, we adopt the velocity-based variant: $\mathbf{Y}^{l+1}=\lambda\bigl[\tau(\mathbf{A}-\mathbf{I})\mathbf{X}^{l}+\mathbf{Y}^{l}\bigr]+(1-\lambda)(\mathbf{A}-\mathbf{I})\mathbf{X}^{l}.$ As a optimizer, we use AdamW loshchilov2018decoupled with learning rate $5.0\times 10^{-5}$, warm-up for first 10 % of steps, then linear decay to step 10k. Also, we evaluate every 5k steps on the validation set, reporting PPL and MLM accuracy.
Table 8: Hyper-parameters for different downstream tasks.
GLUE SQuAD
Batch size 32 16
Weight decay [0.1, 0.01] [0.1, 0.01]
Warmup proportion 0.1 0.1
Learning rate decay linear linear
Training epochs 3 3
Learning rate 0.00005 0.00005
A.1.2 MLM Convergence
As shown in Table 9, adding wave terms yields a mild speedup—Mix (+Full) leads early (e.g., +2.6 at 40k and +1.7 at 60k vs. Diffusion)—while Full Wave largely tracks Diffusion thereafter. In short, wave tends to accelerate convergence slightly, but the residual choice does not substantially alter the overall convergence behavior.
Table 9: MLM accuracy vs. steps.
Model 0 20k 40k 60k 80k 100k
Diffusion 0.0 17.38 32.99 40.88 43.51 44.39
Full Wave 0.0 18.40 33.48 40.80 43.67 44.52
Mix (+Full) 0.0 19.40 35.56 42.60 44.77 45.56
A.2 CV Tasks
A.2.1 Implementation Details
To demonstrate that our proposed approach are beneficial to CV tasks, we choose 12-layer DeiT DeiT as our backbone. Also, for evaluating its consistency with the other remedy to avoid over-smoothing, we combine our approach with existing technique such as FeatScale wang2022antioversmooth . When we use 12-layer DeiT, we follow the same training recipe, hyper-parameters, and data augmentation with DeiT ; wang2022antioversmooth . We implemented DeiT based on Pre-LN Wavy Transformer blocks considering that original DeiT is based on Pre-LN Transformer block. We set $\tau=0.5$, and we select $\mathbf{X}^{l+1}=\boldsymbol{\lambda}\,\mathbf{X}_{\text{wave}}^{l+1}+(1-\boldsymbol{\lambda})\,\mathbf{X}_{\text{diffuse}}^{l+1},$ as the mixing strategy with Full Wave and initialize the gating parameter $\boldsymbol{\lambda}$} as an all-zero vector whose length equals the dimension of the hidden states in the mix variant.
A.2.2 Analysis of Over-smoothing for DeiT-based models
(a) Comparison between cosine similarities with 1-sigma interval shading of the conventional DeiT and the DeiT integrated with Wavy Transformer based on Full Wave variant.
(b) Comparison between cosine similarities with 1-sigma interval shading of the DeiT integrated with Wavy Transformer based on Full Wave variant with and without FeatScale.
Figure 4 plots similarity curves for the standard DeiT, DeiT integrated with Wavy Transformer based on Full Wave variant, and Wavy Transformer + FeatScale variant with 1-sigma interval shading. The results clearly show how Wavy Transformer reshapes feature dynamics relative to the conventional DeiT. Integrating FeatScale slightly reduces over-smoothing overall, although its final-layer cosine similarity remains marginally higher than that of Wavy Transformer alone.
A.2.3 Computation of spectral gap and variance metrics
We complement Figure 4 by detailing how we compute the three diagnostics used in the main text (Table 4). Results for spectral gap, node-feature variance, and inter-class variance are reported in the main paper; the definitions here are for reproducibility.
Spectral Gap
For each image, we average multi-head attention over heads to obtain a row-/right-stochastic matrix $\textbf{A}\in\mathbb{R}^{n\times n}$ of the final layer. Let $\{\lambda_{i}\}_{i=1}^{n}$ be the eigenvalues of A ordered by magnitude. The per-sample spectral gap is
$\delta\;=\;1-|\lambda_{2}(\textbf{A})|,$
computed by taking the second-largest eigenvalue magnitude of A (via an eigen decomposition of the head-averaged A), then averaged over the validation set. Smaller $\delta$ indicates weaker mixing and thus less over-smoothing.
Node-Feature Variance
Let $\textbf{X}\in\mathbb{R}^{n\times d}$ be the token (patch) embeddings from a designated layer at evaluation time. We compute the variance across tokens after the final LN for each channel and then average across channels:
$\mathrm{NodeVar}\;=\;\frac{1}{d}\sum_{j=1}^{d}\mathrm{Var}_{i=1..n}\big(\textbf{X}_{ij}\big).$
We report the mean of this scalar over validation images.
Inter-Class Variance
Let $\mathcal{K}$ be the set of classes. Tokens inherit their image labels; for each $k\in\mathcal{K}$, define the class centroid $\mu_{k}$ as the mean token embedding. We then compute the variance across class centroids and average over channels:
$\mathrm{InterClassVar}=\frac{1}{d}\sum_{j=1}^{d}\mathrm{Var}_{k\in\mathcal{K}}\big(\mu_{k,j}\big).$
This measures feature separability (often correlating with accuracy) and is not used as an over-smoothing diagnostic.
A.2.4 Ablation of FFN for Velocity
Table 10 shows the result of ablation of FFN for velocity. In the Full Wave setting, the FFN for velocity yields a modest but consistent gain (+0.53 Top-1), indicating its effectiveness. However, considering the added computational cost, the momentum-only Light Wave is the more practical choice for cost-sensitive use (see Appendix A.4).
Table 10: Ablation: removing Velocity-FFN (DeiT-Tiny, 150 epochs).
Configuration Top-1 (%) $\Delta$
Full Wave (diff+wave+FFN+LN) 68.92
Without FFN for velocity 68.39 –0.53
A.3 Sparse-Graph Tasks
A.3.1 Implimentation Details
We use the momentum-only Light Wave residual for all sparse-graph experiments; the mixing coefficient $\lambda$ is set as a scalar parameter initialized to 0 at the start of training and optimized jointly with the network. All other settings follow the DIFFormer protocol unless otherwise noted. Also, we adopt the DIFFormer wu2023difformer definition of diffusion mixing $\tau$. This is corresponding to the time interval $\tau^{2}$ in Wavy Transformer. Thus, throughout we control both with a single knob and report $\tau$; see Appendix B for more details.
A.3.2 OGB: full results and $\tau$-robustness
We report complete results across depths and the coefficient $\tau$. At larger $\tau$, diffusive stacks exhibit collapse at depth, whereas the wave residual mitigates it (Table 11).
Table 11: OGB results (3 seeds). Arxiv: accuracy (%); Proteins: ROC-AUC (%).
Dataset Depth Diffusion + Light Wave $\Delta$
Arxiv 3 68.77 $\pm$ 0.28 66.68 $\pm$ 0.66 -2.09
Arxiv 5 61.72 $\pm$ 1.11 67.06 $\pm$ 0.03 +5.34
Arxiv 7 24.44 $\pm$ 4.51 66.73 $\pm$ 0.33 +42.29
Proteins 3 79.26 $\pm$ 0.40 79.58 $\pm$ 0.27 +0.32
Proteins 5 69.42 $\pm$ 2.31 80.14 $\pm$ 0.67 +10.72
A.3.3 Citation Graphs Benchmark
[…]
Dataset Depth Diffusion + Light Wave $\Delta$
Cora 4 80.02 $\pm$ 0.29 78.54 $\pm$ 1.14 –1.48
Cora 16 83.28 $\pm$ 0.76 84.22 $\pm$ 0.94 +0.94
Cora 20 39.92 $\pm$ 5.14 85.18 $\pm$ 0.48 +45.26
Citeseer 4 72.04 $\pm$ 0.25 71.74 $\pm$ 0.27 –0.30
Citeseer 8 62.86 $\pm$ 17.73 69.32 $\pm$ 13.26 +6.46
Citeseer 10 34.76 $\pm$ 1.44 34.92 $\pm$ 2.49 +0.16
PubMed 4 75.74 $\pm$ 0.73 74.76 $\pm$ 1.15 –0.98
[…]
PubMed 20 58.68 $\pm$ 2.73 79.22 $\pm$ 0.93 +20.54
Table 13: Citation graphs at $\tau=0.5$. $\ddagger$: both models collapse.
Dataset Depth Diffusive + Light Wave $\Delta$
Cora 12 29.40 $\pm$ 0.00 60.38 $\pm$ 24.7 +31.0
Cora 20 29.40 $\pm$ 0.00 29.00 $\pm$ 0.00 $\pm$0.0
Citeseer 10 22.94 $\pm$ 0.31 22.84 $\pm$ 2.17 –0.10
PubMed 12 49.16 $\pm$ 9.70 69.74 $\pm$ 15.7 +20.6
PubMed 20 29.70 $\pm$ 0.00 40.10 $\pm$ 0.84 +10.4
A.3.4 Cosine similarity across depth
To quantify over-smoothing for sparse-graph benchmarks, we track the pairwise cosine similarity between node embeddings after each layer in a 20-layer stack on Cora with $\tau=0.2$. We report means over 5 seeds. Higher values (approaching 1.0) indicate stronger representation collapse within a layer. As shown in Table 14, the diffusion-only residual rapidly saturates to $\approx\!1.00$, whereas the wave residual maintains substantially lower similarity across depth, preserving representation diversity.
Table 14: Layer-wise pairwise cosine similarity across node features on Cora (20 layers; 5 seeds). Higher means more over-smoothing.
Model Layer 1 Layer 4 Layer 8 Layer 12 Layer 16 Layer 20
Diffusion 0.99 1.00 1.00 1.00 1.00 1.00
Wave 0.31 0.058 0.087 0.13 0.23 0.38
A.4 Full Results of Computational Cost
[…]
Variant Inference (k tokens/s) Training (ms/iter) Peak GPU (GB)
Diffusion 101.6 415.6 18.31
+ Full Wave 63.3 1031 28.71
+ Light Wave 101.3 436.2 18.69
[…]
Variant Inference (img/s) Training (ms/iter) Peak GPU (GB)
Diffusion 2631.1 618.6 8.25
+ Full Wave 1312.2 1023.4 20.31
+ Light Wave 2644.2 617.6 9.14
A.4.1 Throughput–Accuracy Trade-off Comparison for Full Wave Branch
Table 17 compares the inference throughput (images/s) of various transformer variants, measured on the ImageNet validation set using V100 GPUs and a batch size of 256. Introducing wave dynamics (Diffuse+Wave or Wave alone) reduces throughput by roughly 50% from about 2 600 to 1 200–1 300 images per second which may pose a limitation in real-time applications. However, this slowdown must be balanced against the gains in representational power and accuracy that wave-enhanced models provide. A simple mitigation is to insert wavy blocks in only a subset of layers to recover part of the lost throughput. Concretely, replacing only the last six blocks with wavy blocks (Wave (6)) increases throughput by about 33 % relative to the full Diffuse+Wave with FeatScale variant (1 649 img/s vs. 1 241 img/s) while sacrificing 0.18 pt in Top-1 accuracy (72.44 % vs. 72.62 %). In this experiment, the velocity tensor was initialized to zero at the first wavy block.
Model Residual Connections Throughput Top-1 (%)
DeiT-Ti DeiT Diffusion 2632.1 72.17
DeiT-Ti + FeatScale wang2022antioversmooth Diffusion 2483.2 72.35
DeiT-Ti Diffusion+ Full Wave 1266.0 72.33
DeiT-Ti + FeatScale Diffusion+ Full Wave 1241.4 72.62
DeiT-Ti Full Wave 1312.2
DeiT-Ti + FeatScale Diffuse+ Full Wave (6) 1649.3 72.44
A.5 Wavy Transformer under causal attention
For supplemental information, we test the Light Wave residual under strictly causal attention using nanoGPT nanogpt . We use the OpenWebText dataset and initialize from an OpenAI GPT-2 checkpoint, keeping the standard GPT causal mask and training loop unchanged. During fine-tuning, we use a batch size of 1 with gradient accumulation over 32 steps; training runs for 50 iterations with a learning rate of $3\times 10^{-5}$ and learning-rate decay enabled. Under this setup, the best validation loss, averaged over three seeds, was $3.0178$ for the Light Wave mix (scalar $\lambda$ initialized to 0) versus $3.0196$ for the diffusion baseline—an absolute gap of $0.0018$ ($\approx 0.06\%$); while the Light Wave mix is slightly better, we do not observe a meaningful difference in validation loss. A more comprehensive assessment under causal masking requires substantially longer training; we therefore view this as worthwhile future work, related to potential over-smoothing in causal-attention Transformers. From a diffusive-dynamics perspective, the attention matrix acts as a diffusion-coefficient matrix; under causal masking it becomes strictly unidirectional along the sequence, so causal attention can be viewed as diffusion on a directed complete graph with unidirectional edges.
Appendix B Implementation Details of Light Wave
As shown in Section 3.2, if we directly discretize Eq. (9) in time, rather than bypassing it through Eq. (10), we obtain
${\bf X}^{l+1}=\tau^{2}{\bf A}{\bf X}^{l}+(1-\tau^{2}){\bf X}^{l}+({\bf X}^{l}-{\bf X}^{l-1}).$ (16)
This equation can be seen as adding the momentum term $({\bf X}^{l}-{\bf X}^{l-1})$ into the usual diffusive attention without considering the coefficient $\tau^{2}$. Also, if we consider to blend the outputs of the diffusive attention layer and the proposed wavy attention layer:
$\mathbf{X}^{l+1}=\boldsymbol{\lambda}\,\mathbf{X}_{\text{wave}}^{l+1}+(1-\boldsymbol{\lambda})\,\mathbf{X}_{\text{diffuse}}^{l+1},$
where the mixing vector $\boldsymbol{\lambda}\in[0,1]^{d}$ is defined as $\boldsymbol{\lambda}=\textrm{sigmoid}(\boldsymbol{\theta}),$ with $\boldsymbol{\theta}\in\mathbb{R}^{d}$ a trainable parameter vector.
Decoupling $\tau$ Across Diffusion and Wave.
Let $\tau_{\mathrm{d}}$ and $\tau_{\mathrm{w}}$ be the (possibly different) time intervals used in the diffusive and wavy paths, respectively. Substituting
$\mathbf{X}_{\text{diffuse}}^{l+1}$ $=(1-\tau_{\mathrm{d}})\,\mathbf{X}^{l}+\tau_{\mathrm{d}}\,\mathbf{A}\mathbf{X}^{l},$ (17)
$\mathbf{X}_{\text{wave}}^{l+1}$ $=(1-\tau_{\mathrm{w}}^{2})\,\mathbf{X}^{l}+\tau_{\mathrm{w}}^{2}\,\mathbf{A}\mathbf{X}^{l}+\bigl(\mathbf{X}^{l}-\mathbf{X}^{l-1}\bigr),$ (18)
into the blend gives the elementwise form
$\mathbf{X}^{l+1}\;=\;(1-\bar{\rho})\,\mathbf{X}^{l}+\bar{\rho}\,\mathbf{A}\mathbf{X}^{l}\;+\;\boldsymbol{\lambda}\bigl(\mathbf{X}^{l}-\mathbf{X}^{l-1}\bigr),\qquad\bar{\rho}\;:=\;\boldsymbol{\lambda}\tau_{\mathrm{w}}^{2}+(1-\boldsymbol{\lambda})\tau_{\mathrm{d}}.$ (19)
Equation (19) shows that it is not necessary to strictly match $\tau$ between the two paths: only the effective diffusion mix $\bar{\rho}$ and the momentum weight $\boldsymbol{\lambda}$ govern the update. Equivalently, by reparameterizing with a single knob $\tau:=\bar{\rho}\in[0,1]$, the update reduces to the Light Wave form used in our experiments.
Appendix C Energy Evolution in Diffusion and Wave Systems